OpenStudy (anonymous):

lim x→−7 of 7 − |x| / 7 + x

OpenStudy (tkhunny):

Do you mean \(\dfrac{7-|x|}{7+x}\), because that is very much NOT what you have written. You have \(7 - \dfrac{|x|}{7} + x\). Remember your Order of Operations.

OpenStudy (anonymous):

yes i meant what you think i meant but i had a hard time typing sorry

OpenStudy (anonymous):

limit((7-abs(x))/(7+x), x = -7) = 1

OpenStudy (anonymous):

tkhunny can you please help me

OpenStudy (tkhunny):

Kind of a fun little idea on this one. We have the usual... If x >= 0, then |x| = x If x < 0, then |x| = -x However, Evaluating the limit, eventually we get nowhere near x = 0 and we can just ignore the part about x >= 0. Considering, then, ONLY x < 0, we have \(\lim_{x \rightarrow -7}\dfrac{7 - |x|}{7+x} = \lim_{x \rightarrow -7}\dfrac{7 - (-x)}{7+x}\) and the whole thing is a lot simpler.

OpenStudy (anonymous):

wait... i thought that the absolute value of ANY number positive or negative is always positive

OpenStudy (anonymous):

the absolutely value of -2 is 2

OpenStudy (anonymous):

if what you are saying is correct then the evaluated number would be 0 ???

OpenStudy (tkhunny):

|-2| = -(-2) = +2 If x < 0, |x| = -x

OpenStudy (anonymous):

im still confused

OpenStudy (tkhunny):

If x > 0, |x| = x Example |5| = 5 If x < 0, |x| = -x Example |-3| = -(-3) = +3 You have to get his idea functionally in your head. Essentially, the absolute value does nothing to positive numbers. The absolute value changes the sign on negative numbers.

OpenStudy (anonymous):

ok i understand.. but i still dont understand my initial problem

OpenStudy (tkhunny):

\(\dfrac{7-(-x)}{7+x} = \dfrac{7+x}{7+x} = 1,\;except\;where\;x = -7.\) It's not continuous, but the limit exists.